The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 68 January 22nd, 2003

Contents

Letter from the Editor  1
Abstracts of 12 Refereed Papers 2
Recent astro-ph Listings  9

Letter From The Editor

Dear Colleagues,

On behalf of the Magellanic Clouds Working Group, I’d like to extend our best wishes to you in the new year. May your research be fruitful, your proposals be accepted, and your next trip around the Sun be a good one! The Magellanic Clouds Working Group is pleased to bring you issue 68 of the Magellanic Clouds Newsletter. This month, we present 12 submitted abstracts and a collection of recent astro-ph listings of interest to Magellanic Clouds researchers. As always, MC News is available in a variety of formats, including L^{A}TEX, PostScript, Adobe Acrobat (PDF), and HTML. To download your preferred format, please direct your web browser to http://www.astro.uiuc.edu/projects/mcnews/MCNews.shtml.

Best Wishes, Bryan Dunne Editor, MC News

Abstracts of Refereed Papers

The Gaseous Disk Almost Perpendicular to the Galactic Plane

N. V. Bystrova

St. Petersburg Branch of the Special Astrophysical Observatory RAS

Around the calculated position of the Magellanic Stream’s great circle two very extended zones (I and II) were chosen situated azimuthally opposed each other. These zones were observed with the RATAN-600 telescope in the HI radio-line at the relative low LSR-velocities. Six hours were necessary to cross the zones at every declination from −40 to +62.5 degrees. The comparison was made of the warp-like signals in the Magellanic Stream and those in the Galactic gaseous disk (the North and the South warps). Their sizes, the amount and the direction of the displacements relative to the RA direction in the Magellanic Stream were evaluated. The position of the plane through maximum deformations was determined. The behaviour of HI in the Stream shows that there is possibly a gaseous disk similar to that of the polar ring in external galaxy. The warp tendencies in the Stream’s plane were firstly proposed using the observations with the Large Pulkovo Radio Telescope.

Comments: To appear in Astronomical and Astrophysical Transactions 2002. e-mail: bnv@fsao.spb.su

Results of the ESO-SEST Key Programme on CO in the Magellanic Clouds. IX. The giant LMC HII region complex N11

F.P. Israel (1), Th. de Graauw (2), L.E.B. Johansson (3), R.S. Booth (3) F. Boulanger (4), G. Garay (5), M.L. Kutner (6), J. Lequeux (7), L-A. Nyman (8), M. Rubio (5)

(1) Sterrewacht Leiden, Netherlands  (2) SRON Groningen, Netherlands
(3) OSO Onsala, Sweden  (4) ENS and IAS, Paris, France
(5) Univ. de Chile Santiago, Chile  (6) University of Texas, USA
(7) Obs. de Paris, France  (8) ESO Santiago, Chile

We present maps and a catalogue containing the J=1-0 12CO parameters of 29 individual molecular clouds in the second-brightest LMC star formation complex, N11. In the southwestern part of N11, molecular clouds occur in a ring or shell surrounding the major OB star association LH9. In the northeastern part, a chain of molecular clouds delineates the rim of one of the so-called supergiant shells in the LMC. The well-defined clouds have dimensions close to those of the survey beam (diameters of 25 pc or less). Some of the clouds were also observed in J=2-1 12CO, and in the lower two transitions of 13CO. Clouds mapped with a twice higher angular resolution in J=2-1 12CO show substructure with dimensions once again comparable to those of the mapping beam. The few clouds for which we could model physical parameters have fairly warm (T(kin) = 60 - 150 K) and moderately dense (n(H2) = 3000 cm-3) gas. The northeastern chain of CO clouds, although lacking in diffuse intercloud emission, is characteristic of the more quiescent regions of the LMC and appears to have been subject to relatively little photo-processing. The clouds forming part of the southwestern shell or ring, however, are almost devoid of diffuse intercloud emission and also exhibit other characteristics of an extreme photon-dominated region (PDR).

Comments: A&A, in press WWW: http://xxx.lanl.gov/abs/astro-ph/0211450 e-mail: israel@strw.leidenuniv.nl

The Wind of the B[e] Supergiant Hen S22 Viewed through a Reflection Nebula in DEM L 106

Y.-H. Chu (1), C.-H. R. Chen (1), C. Danforth (2), B. C. Dunne (1), R. A. Gruendl (1), Y. Nazé (3), M. S. Oey (4), S. D. Points (5)

(1) University of Illinois at Urbana-Champaign  (2) Johns Hopkins University
(3) Institut d’Astrophysique et de Géophysique  (4) Lowell Observatory

(5) Northwestern University

Narrow-band HST WFPC2 images reveal a bow-shock-like halo around the H ii region N30B toward the B[e] supergiant Hen S22 located within the larger DEM L 106 nebula in the Large Magellanic Cloud. High-dispersion spectra of N30B show a narrow Hα emission component from the ionized gas; the velocity variations indicate a gas flow of −5 to −10 km/s in the vicinity of the H ii regions, which is resultant from interactions with Hen S22’s stellar wind and responsible for the bow-shock morphology. Spectra of N30B’s halo show broad Hα profiles extending over > 1000 km/s, similar to that of Hen S22, indicating that the halo is a reflection nebula of Hen S22. Broad-band morphologies of N30B’s halo are also consistent with the reflection nebula interpretation. We use dust-scattering properties and the observed brightnesses of the reflection nebula and Hen S22 to constrain the reflection geometry. The reflected stellar Hα emission and absorption vary across the reflection nebula as a result of viewing S22’s anisotropic wind across varying angles. This reflection nebula, together with the edge-on orientation of Hen S22’s disk, provides an invaluable opportunity to study the disk and polar winds of a B[e] supergiant.

Comments: To appear in the Astronomical Journal, April 2003 WWW: http://arXiv.org/abs/astro-ph/0212462 e-mail: chu@astro.uiuc.edu

An Outflow from the Nebula around the LBV Candidate S119

Kerstin Weis (1,2), Wolfgang J. Duschl (1,2), Dominik J. Bomans (3)

(1) Institut für Theoretische Astrophysik, University of Heidelberg, Germany

(2) Max-Planck-Institut für Radioastronomie, Bonn, Germany

(3) Astronomical Institute, Ruhr-Universität Bochum, Germany

We present an analysis of the kinematic and morphological structure of the nebula around the LMC LBV candidate S 119. On HST images, we find a predominantly spherical nebula which, however, seems to be much better confined in its eastern hemisphere than in the western one. The filamentary western part of the nebula is indicative of matter flowing out of the nebula’s main body. This outflow is even more evidenced by our long-slit echelle spectra. They show that, while most of the nebula has an expansion velocity of 25.5 km/s, the outflowing material reaches velocities of almost 140 km/s, relative to the systemic one. A ROSAT HRI image shows no trace of S 119 and thus no indications of hot or shocked material.

Comments: Accepted by A&A WWW: http://www.mpifr-bonn.mpg.de/staff/kweis/publications.html e-mail: kweis@mpifr-bonn.mpg.de

X-Ray Emission from Multi-Phase Shock in the Large Magellanic Cloud Supernova Remnant N49

Sangwook Park (1), David N. Burrows (1), Gordon P. Garmire (1), John A. Nousek (1), John P. Hughes (2), and Rosa Murphy Williams (3)

(1) Department of Astronomy and Astrophysics, Penn State University

(2) Department of Physics and Astronomy, Rutgers University

(3) Department of Astronomy, University of Massachusetts

The supernova remnant (SNR) N49 in the Large Magellanic Cloud has been observed with the Advanced CCD Imaging Spectrometer (ACIS) on board the Chandra X-Ray Observatory. The superb angular resolution of the Chandra/ACIS images resolves a point source, the likely X-ray counterpart of soft gamma-ray repeater SGR 0526−66, and the diffuse filaments and knots across the SNR. These filamentary features represent the blast wave sweeping through the ambient interstellar medium and nearby dense molecular clouds. We detect metal-rich ejecta beyond the main blast wave shock boundary in the southwest of the SNR, which appear to be explosion fragments or “bullets” ejected from the progenitor star. The detection of strong H-like Si line emission in the eastern side of the SNR requires multi-phase shocks in order to describe the observed X-ray spectrum, whereas such a multi-phase plasma is not evident in the western side. This complex spectral structure of N49 suggests that the post-shock regions toward the east of the SNR might have been reheated by the reverse shock off the dense molecular clouds while the blast wave shock front has decelerated as it propagates into the dense clouds. The X-ray spectrum of the detected point-like source is continuum dominated and can be described with a power law of Γ ∼ 3. This provides a confirmation that this point-like X-ray source is the counterpart of SGR 0526−66 in the quiescent state.

Comments: Accepted by ApJ (v586, March 20, 2003) WWW: http://xxx.lanl.gov/abs/astro-ph/0211552 e-mail: park@astro.psu.edu

Constraints on the Ionization Balance of Hot-Star Winds from FUSE Observations of O Stars in the Large Magellanic Cloud

D. Massa (1), A.W. Fullerton (2), G. Sonneborn (3), J.B. Hutchings (4)

(1) SGT Inc.  (2) Johns Hopkins University
(3) NASA’s GSFC  (4) HIA, NRC of Canada

We use a Sobolev with Exact Integration model to analyze the winds lines of 25 LMC O stars. The data include FUSE profiles of C III, N III, S IV, P V, S VI, and O VI and IUE or HST data for Si IV, C IV, and N V. Several of the FUSE lines are unsaturated, so meaningful optical depths (equivalently, mass loss rate times ionization fractions), as a function of wind velocity can be determined. Ratios of these quantities give the relative ionization as a function of velocity and demonstrate that, except for O VI in all stars and S VI in the later stars, the wind ionization shifts toward lower stages at higher velocity. Because O VI and S VI do not behave like the other ions, they must be produced by a different mechanism. Using mass-loss rates determined from the Vink et al. relationships, we derive mean ionization fractions. Because these are all less than one, the derived mass loss rates cannot be too small. However, the ion fractions for P V (expected to be dominant in some winds), never exceed 0.20. This implies that either the calculated mass loss rates or the assumed P abundances are too large, or the winds are strongly clumped. We examine correlations between the mean ion fractions and stellar parameters, and find two significant relationships. First, as expected, the mean ionization fraction of lower ions decreases with increasing temperature. Second, the mean ionization fraction of S VI in the latest stars and O VI in all stars increases with terminal velocity, re-affirming Cassinelli and Olson’s conjecture that O VI is produced non-radiatively. Finally, we discuss peculiar aspects of three stars, BI 272, BI 208, and Sk-67 166.

Comments: To appear in ApJ, April 1, 2003 WWW: http://xxx.lanl.gov/abs/astro-ph/0211518 e-mail: massa@taotaomona.gsfc.nasa.gov

An X-ray investigation of the NGC 346 field in the SMC (2) : the field population

Y. Nazé (1,2), J.M. Hartwell (3), I.R. Stevens (3), J. Manfroid (1,4,5), S. Marchenko (6), M. F. Corcoran (7), A.F.J. Moffat (8), G. Skalkowski (8)

(1) Institut d’Astrophysique et de Géophysique, Université de Liège, Allée du 6 Août 17, Bat. B5c, B 4000 - Liège (Belgium)

(2) Research Fellow F.N.R.S.

(3) School of Physics & Astronomy, University of Birmingham, Edgbaston, Birmingham B15 2TT (UK)

(4) Research Director F.N.R.S.

(5) Visiting Astronomer, European Southern Observatory

(6) Department of Physics and Astronomy, Thompson Complex Central Wing, Western Kentucky University, Bowling Green, KY 42101-3576 (USA)

(7) Universities Space Research Association, High Energy Astrophysics Science Archive Research Center, Goddard Space Flight Center, Greenbelt, MD 20771

(8) Département de physique, Université de Montreal, C.P. 6128, Succ. Centre-Ville, Montreal, QC, H3C 3J7 (Canada)

We present results from a Chandra observation of the NGC 346 cluster, which is the ionizing source of N66, the most luminous H ii region and the largest star formation region in the SMC. In the first part of this investigation, we have analysed the X-ray properties of the cluster itself and the remarkable star HD 5980. But the field contains additional objects of interest. In total, 75 X-ray point sources were detected in the Chandra observation: this is five times the number of sources detected by previous X-ray surveys. We investigate here their characteristics in detail. Due to high foreground absorption, the sources possess rather high hardness ratios. Their cumulative luminosity function appears generally steeper than that for the rest of the SMC at higher luminosities. Their absorption columns suggest that most of the sources belong to NGC 346. Using DSS data and new UBV RI imaging with the ESO 2.2m telescope, we also discovered possible counterparts for 32 of these X-ray sources and estimated a B spectral type for a large number of these counterparts. This tends to suggest that most of the X-ray sources in the field are in fact X-ray binaries. Finally, some objects show X-ray and/or optical variability, with a need for further monitoring.

Comments: Accepted by ApJ, also on the web as astro-ph/0212197 WWW: http://vela.astro.ulg.ac.be/Preprints/P79/index.html e-mail: naze@astro.ulg.ac.be

Active star formation in N11B Nebula in the Large Magellanic Cloud: a sequential star formation scenario confirmed

Rodolfo H. Barb(1), Monica Rubio(2), Miguel R. Roth(3), Jorge Garca(4)

(1) Facultad de Ciencias Astronmicas y Geofsicas, Universidad Nacional de La Plata, La Plata, Argentina and Departamento de Fsica, Universidad de La Serena, Chile

(2) Departamento de Astronoma, Universidad de Chile, Santiago, Chile

(3) Las Campanas Observatory, The Observatories, Carnegie Institution of Washington, La Serena, Chile

(4) Gemini Observatory Southern Operations Center, La Serena, Chile

The second largest HII region in the Large Magellanic Cloud, N11B has been surveyed in the near IR.We present JHKs images of the N11B nebula.These images are combined with CO(1-0) emission line data and with archival NTT and HST/WFPC2 optical images to address the star formation activity of the region.IR photometry of all the IR sources detected is given.We confirm that a second generation of stars is currently forming in the N11B region. Our IR images show the presence of several bright IR sources which appear located towards the molecular cloud as seen from the CO emission in the area.Several of these sources show IR colours with YSO characteristics and they are prime candidates to be intermediate-mass Herbig Ae/Be stars.For the first time an extragalactic methanol maser is directly associated with IR sources embedded in a molecular core.Two IR sources are found at 2”(0.5 pc) of the methanol maser reported position.Additionally, we present the association of the N11A compact HII region to the molecular gas where we find that the young massive O stars have eroded a cavity in the parental molecular cloud, typical of a champagne flow. The N11 region turns out to be a very good laboratory for studying the interaction of winds, UV radiation and molecular gas. Several photo-dissociation regions are found.

Comments: 11 figures, accepted in Astronomical Journal (April 2003) WWW: http://arxiv.org/abs/astro-ph/0212525 e-mail: rbarba@dfuls.cl

Tight LMC massive star clusters R 127 and R 128

M. Heydari-Malayeri (1), F. Meynadier (1) and Nolan R. Walborn (2)

(1) LERMA, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France

(2) Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, Maryland 21218, USA

We study the Large Magellanic Cloud (LMC) star clusters R 127 and R 128 using imaging and spectroscopy obtained at the ESO NTT telescope. An advanced image restoration technique allows us to resolve these two clusters into at least 14 and 33 stars respectively and obtain their photometry. In particular, we show that the core of R 127 is composed of at least four stars and identify the Luminous Blue Variable (LBV) component. The closest neighbor of the LBV (star #8) is 1′′.5 away. Moreover, from medium dispersion spectroscopy we determine the spectral types for 19 stars in and near both clusters, and in particular present the first spatially resolved observation of the second brightest component of the R 127 cluster (star #3) situated 3′′.3 from the LBV. By comparing with evolutionary models we also look into the stellar ages. The oldest stars of the cluster are ∼ 6–8 Myr old, whereas the most massive star of the region (#7), formed ∼ 3 Myr ago as an 80 M_{⊙} star, has turned into an LBV, the “R 127” star.

Comments: Accepted by A&A WWW: http://fr.arxiv.org/abs/astro-ph/0301298 e-mail: Frederic.Meynadier@obspm.fr

Ages and Metallicities of Eight Star Clusters and Their Surrounding Fields in the Inner Disk of the Large Magellanic Cloud

D. Geisler (1), A. E. Piatti (2), E. Bica (3), J. J. Claria (4)

(1) Universidad de Concepcion, Concepcion, Chile

(2) Instituto de Astronomia y Fisica del Espacio, B.A., Arg.

(3) Universidade Federal do Rio Grande do Sul, P.A., Brazil

(4) Observatorio Astronomico, Cordoba, Argentina

We present Washington system CMDs for 8 star clusters and their surrounding fields which, with one exception, lie within the inner parts of the LMC disk. Careful attention is paid to separating out the cluster and field star distributions. Ages and metallicities are then determined in a consistent manner for both populations in two different ways. All clusters are found to be of intermediate age (1-3 Gyr), with [Fe/H] from -0.4 to -0.9. We find that the stellar population of each cluster is similar to that of the field where it is embedded, sharing its mean age and metallicity. Combining the present sample with a revision of that of Bica et al. (1998) studied similarly, we find that our metallicities for intermediate age clusters are intermediate to those for clusters of similar age studied by Olszewski et al. (1991) and by Beasley, Hoyle and Sharples (2002). A combined age-metallicity relation is presented which shows that LMC clusters formed between 1-3 Gyr ago with a mean metallicity (-0.5 dex) and metallicity spread (0.23 dex) independent of age. Good agreement is found with the bursting model of Pagel and Tautvaisiene (1998). No evidence for a metallicity gradient is found.

e-mail: doug@kukita.cfm.udec.cl

Long Period Variables detected by ISO in the SMC

M.-R.L. Cioni (1), J.A.D.L. Blommaert (2), M.A.T. Groenewegen (2), H.J. Habing (3),

J. Hron (4), F. Kerschbaum (4), C. Loup (5), A. Omont (5), J.Th. van Loon (9), L.B.F.M. Waters (2,6), P.A. Whitelock (7), A. Zijlstra (8)

(1) European Southern Observatory, Karl–Schwarzschild–Str. 2, D–85748 Garching bei München Germany

(2) Institute of Astronomy, Katholieke Universiteit Leuven, Celestijnenlaan 200B, B3001, Leuven, Belgium

(3) Sterrewacht Leiden, Niels Bohrweg 2, 2333 RA Leiden, The Netherlands

(3) Institut für Astronomie der Universität Wien, Türkenschanzstrasse 17, 1180 Wien, Austria

(4) Institute d’Astrophysique de Paris, CNRS, 98bis Bld. Arago. 75014 Paris, France

(5) Astronomical Institute “Anton Pannekoek”, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

(6) South African Astronomical Observatory, P.O.Box 9, 7935 Observatory, Republic of South Africa

(7) Department of Physics, UMIST, P.O. Box 88, Manchester M60 1QD, UK

(8) Astrophysics Group, School of Chemistry and Physics, Keele University, Staffordshire ST5 5BG, UK

This article presents the study of the light–curves extracted from the MACHO database of a sample of stars observed by the Infrared Space Observatory in the Small Magellanic Cloud. These stars belong to the ISO–Mini-Survey catalogue of the Magellanic Clouds (ISO–MCMS, Loup et al.). Most of them are in the asymptotic giant branch (AGB) and supergiant phases. The dominant period and amplitude of pulsation have been derived and the stars have been classified as Mira or Semi–Regular pulsators. Furthermore, the cross–identification with near–infrared DENIS and 2MASS magnitudes available within the ISO–MCMS allowed us: (i) to investigate the properties of these variables in the combined near– and mid–infrared colour–magnitude diagrams, (ii) to derive the bolometric magnitude by integrating the spectral energy distribution and (iii) to calculate the mass–loss rate. The stars have been divided into carbon– (C–) and oxygen–rich (O–rich) using the (J − K_{S}, K_{S}) colour–magnitude diagram and their period and amplitude distributions have been compared. C–rich AGB stars dominate at periods of about 250 days and have on average a larger amplitude than O–rich AGB stars. This effect, not previously detected from the study of similar stars in the Large Magellanic Cloud and in the Baade’s window, might be closely related to the metallicity of the environment in which the stars have formed.

Comments: 14 pages, 10 figures, submitted to A&A e-mail: mcioni@eso.org

Fourier decomposition and frequency analysis of the pulsating stars

with P<1 d in the OGLE database II. Multiperiodic RR Lyr variables in the Galactic Bulge

Pawel Moskalik (1) & Ennio Poretti (2)

(1) Copernicus Astronomical Center, Warsaw, Poland

(2) INAF-Osservatorio di Brera, Merate, Italy

We present the results of a systematic search for multiperiodic pulsators among the Galactic Bulge RR Lyrae stars of the OGLE-1 sample. We identify one ”canonical” double-mode variable (RRd star) pulsating in two radial modes. In 38 stars we detect secondary periodicities very close to the primary pulsation frequency. This type of multiperiodic variables constitute 23RRc population of the Bulge. With the observed period ratios of 0.95-1.02 the secondary periods must correspond to nonradial modes of oscillation. Their beating with the primary (radial) pulsation leads to a long-term amplitude and phase modulation, known as the Blazhko effect. The Blazhko RRab variables occur more frequently in the Galactic Bulge than in the LMC. The opposite tendency is seen in case of the RRd stars. The differences of incidence rates are most likely caused by different metallicity of the two populations. We discuss pulsation properties of the OGLE-1 Blazhko stars and compare them with predictions of theoretical models. We argue, that the oblique magnetic pulsator model of Shibahashi (2000) cannot account for the observations and should be ruled out.

Comments: Accepted for publication in Astronomy & Astrophysics WWW: http://arXiv.org/abs/astro-ph/0211296 e-mail: poretti@merate.mi.astro.it

Recent astro-ph Listings

astro-ph/0301298:
Title: Tight LMC massive star clusters R 127 and R 128
Authors: M. Heydari-Malayeri, F. Meynadier, Nolan R. Walborn
Comments: 16 pages, 9 figures, accepted for publication in A&A
astro-ph/0301296:
Title: Fundamental Properties and Distances of LMC Eclipsing Binaries: IV. HV 5936
Authors: E.L. Fitzpatrick, I. Ribas, E.F. Guinan, F.P. Maloney, A. Claret
Comments: 44 pages, including 8 pages of tables and 11 pages of figures.
Accepted for publication in the April 20, 2003 issue of the Astrophysical Journal
astro-ph/0301291:
Title: Calibration of the Distance Scale from Cepheids
Authors: Pascal Fouque, Jesper Storm, Wolfgang Gieren
Comments: to be published in: "Stellar Candles", Lecture Notes in Physics (this http URL)
astro-ph/0301279:
Title: Distances from Planetary Nebulae
Authors: Robin Ciardullo
Comments: Invited review at the International Workshop on "Stellar Candles for the
Extragalactic Distance Scale", held in Concepcion, Chile
astro-ph/0301208:
Title: Spectral variations of the X-ray binary pulsar LMC X-4 during its long period
intensity variation and a comparison with Her X-1
Authors: S. Naik, B. Paul
Comments: 7 pages, 5 figures, Accepted for publication in Astronomy and Astrophysics
astro-ph/0301195:
Title: Hard X-ray timing and spectral properties of PSR B0540-69
Authors: J. de Plaa, L. Kuiper, W. Hermsen
Comments: 7 pages and 4 figures. Accepted for publication in Astronomy & Astrophysics
astro-ph/0301100:
Title: Current Uncertainties in the Use of Cepheids as Distance Indicators

Author: Michael Feast
Comments: 26 pages, no figs, to appear in "Stellar Candles for the Extragalactic
Distance Scale" Lecture Notes in Physics (Springer) ed. D. Alloin and W. Gieren
astro-ph/0301055:
Title: Three New Long Period X-ray Pulsars Discovered in the Small Magellanic Cloud
Authors: D.J. Macomb, D.W. Fox, R.C. Lamb, T.A. Prince
Comments: Accepted for publication in the Astrophysical Journal
astro-ph/0212525:
Title: Active star formation in N11B Nebula in the Large Magellanic Cloud:
A sequential star formation scenario confirmed
Authors: Rodolfo H. Barba, Monica Rubio, Miguel R. Roth, Jorge Garcia
Comments: 11 figures, accepted in Astronomical Journal (April 2003), needs AASTEX package
astro-ph/0212498:
Title: XMM-Newton Studies of the Source Population and the Hot Interstellar Medium
in Nearby Galaxies
Authors: W. Pietsch
Comments: 8 pages, 10 figures, uses VisionStyle.cls, Proc. Symnposium ``New Visions of the
X-ray Universe in the XMM-Newton and Chandra Era", 26-30 November 2001, ESTEC,
The Netherlands, a version with full resolution images is available at this http URL
astro-ph/0212462:
Title: The Wind of the B[e] Supergiant Hen S22 Viewed through a Reflection Nebula
in DEM L 106
Authors: Y.-H. Chu, C.-H. R. Chen, C. Danforth, B. C. Dunne, R. A. Gruendl,
Y. Naze, M. S. Oey, S. D. Points
Comments: 19 pages, 7 figures, Astronomical Journal, accepted for April 2003 issue
astro-ph/0212410:
Title: The Optical Gravitational Lensing Experiment. Catalog of RR Lyrae Stars
from the Small Magellanic Cloud
Authors: I. Soszynski, A. Udalski, M. Szymanski, M. Kubiak, G. Pietrzynski,
P. Wozniak, K. Zebrun, O. Szewczyk, L. Wyrzykowski
Comments: 28 pages, 6 figures, LaTex
Journal-ref: Acta Astronomica (2002) 52, 369
astro-ph/0212361:
Title: Non-LTE line-formation for CNO
Authors: Norbert Przybilla, Rolf-Peter Kudritzki, Keith Butler, Sylvia R. Becker
Comments: 6 pages, 2 figures; to be published in "CNO in the Universe",
ASP Conf. Series, eds. C. Charbonnel, D. Schaerer & G. Meynet
astro-ph/0212354:
Title: Far-Ultraviolet & Halpha Spectroscopy of SNR0057-7226 in
the SMC HII Region N66
Authors: Charles Danforth, Ravi Sankrit, William Blair, Chris Howk, You-Hua Chu
Comments: 15 pages in emulateApJ format. Scheduled for April 1, 2003, ApJ
astro-ph/0212319:
Title: Deep XMM-Newton observation of a northern LMC field:
I. Selected X-ray sources
Authors: F. Haberl, K. Dennerl, W. Pietsch (MPE, Garching)
Comments: 11 pages, LaTex, 6 figures, submitted to A&A
astro-ph/0212258:
Title: New Nitrogen and Carbon in AF-supergiants
Authors: Kim A. Venn, Norbert Przybilla

Comments: 9 pages, conference proceeding in CNO in the Universe, ASP Conf. Series,
editted by C. Charbonnel, D. Schaerer, and G. Meynet
astro-ph/0212213:
Title: The Evolution Of LMC X-4 Flares: Evidence For Super-Eddington Radiation
Oozing Through Inhomogeneous Polar Cap Accretion Flows ?
Authors: Dae-Sik Moon, Stephen S. Eikenberry, Ira M. Wasserman (Cornell University)
Comments: 39 pages (including figures and tables), accepted for publication in ApJ
astro-ph/0212212:
Title: Double-periodic blue variables in the Magellanic Clouds
Authors: R.E. Mennickent, G. Pietrzynski, M. Diaz, W. Gieren
Comments: submitted to A&A letters
astro-ph/0212197:
Title: An X-ray investigation of the NGC 346 field in the SMC (2) : the field population
Authors: Y. Naze, J.M. Hartwell, I.R. Stevens, J. Manfroid, S. Marchenko,
M. F. Corcoran, A.F.J. Moffat, G. Skalkowski
Comments: 20 pages, 10 figures (3 in jpg), accepted by ApJ
astro-ph/0212144:
Title: Massive Variability Search and Monitoring by OGLE and ASAS
Authors: Bohdan Paczynski
Comments: 4 pages, latex, to appear in: ``Towards an International Virtual
Observatory'', June 2002, Garching bei Muenchen (Germany), eds. Gorski K. M. et al,
ESO conference series
astro-ph/0212118:
Title: Iron-Rich Ejecta in the Supernova Remnant DEM L71
Authors: John P. Hughes, Parviz Ghavamian, Cara E. Rakowski, Patrick O. Slane
Comments: 5 pages, including 3 postscript figs, LaTeX, to appear in ApJ Letters
astro-ph/0212106:
Title: Reconstructing Cepheid Light Curve with Fourier Techniques I:
The Fourier Expansion and Interrelations
Authors: Chow-Choong Ngeow, Shashi M. Kanbur, Sergei Nikolaev, Nial R. Tanvir,
Martin A. Hendry
Comments: LaTex, 52 pages, 8 tables, 24 figures, some figures are in .jpg,
ApJ preprint style, ApJ accepted (April 2003)
astro-ph/0212088:
Title: The Opacity of Nearby Galaxies from Counts of Background Galaxies: II.
Limits of the Synthetic Field Method
Authors: Rosa A. Gonzalez, Laurent Loinard, Ronald J. Allen, Sebastien Muller
Comments: 45 pages, 10 Postscript figures
astro-ph/0211558:
Title: Spectral properties of the X-ray binary pulsar LMC X-4 during different
intensity states
Authors: S. Naik, B. Paul
Comments: 4 pages, 4 figures
Journal-ref: Jour. of Astrophysics and Astronomy (2002), 23, 27-31
astro-ph/0211552:
Title: X-Ray Emission from Multi-Phase Shock in the Large Magellanic Cloud
Supernova Remnant N49
Authors: Sangwook Park, David N. Burrows, Gordon P. Garmire, John A. Nousek,
John P. Hughes, Rosa Murphy Williams
Comments: Total 22 pages (in ApJ emulater style), including 8 figures (2 color figs).

Accepted by ApJ For high quality images, contact with park@astro.psu.edu
astro-ph/0211518:
Title: Constraints on the Ionization Balance of Hot-Star Winds from FUSE Observations
of O Stars in the Large Magellanic Cloud
Authors: D. Massa, A.W. Fullerton, G. Sonneborn, J.B. Hutchings
Comments: Accepted for publication in ApJ, 67 pages 8 figures
astro-ph/0211491:
Title: A Supernova Remnant Collision with a Stellar Wind
Authors: Pablo F. Velazquez, Gloria Koenigsberger, Alejandro C. Raga
Comments: 9 files into a tar.gz file (the Latex file of manuscript, 7 jpg figures,
1 style file: aastex.cls). Accepted by ApJ, Vol. 584, 2003 February 10
astro-ph/0211450:
Title: Results of the ESO-SEST Key Programme on CO in the Magellanic Clouds. IX.
The giant LMC HII region complex N11
Authors: F.P. Israel, Th. de Graauw, L.E.B. Johansson, R.S. Booth, F. Boulanger,
G. Garay, M.L. Kutner, J. Lequeux, L.-A. Nyman, M. Rubio
Comments: 14 pages; accepted for publication in A&A
astro-ph/0211296:
Title: Fourier decomposition and frequency analysis of the pulsating stars with P < 1 day
in the OGLE database. II. Multiperiodic RR Lyrae variables in the Galactic Bulge
Authors: P. Moskalik, E. Poretti
Comments: 11 pages, 8 figures, accepted for publication in Astronomy & Astrophysics